OpenStudy (anonymous):

Multiply: 2√11•8√22? 8√11 10√33 64√22 176√2

OpenStudy (unklerhaukus):

\[ 2√11\cdot8√22=2\times8\sqrt{11\times22}\]

OpenStudy (anonymous):

can you simplify 16√222?

OpenStudy (unklerhaukus):

you mean can you simplify \[16\sqrt{242}\]

OpenStudy (anonymous):

11 times 22 is 222?

OpenStudy (unklerhaukus):

\[11\times22=10\times22+1\times22=220+22=242\]

OpenStudy (unklerhaukus):

\[242=2\times 121=2\times11^2\]

OpenStudy (unklerhaukus):

\[2√11\cdot8√22=16\sqrt{2\times11^2}\]

OpenStudy (unklerhaukus):

16x11=176

OpenStudy (anonymous):

i have no idea what youre doing

OpenStudy (unklerhaukus):

no more clues

OpenStudy (anonymous):

You have to \[176\sqrt[]{2}\]